If 3/4 is added to three times a number, the result is 1/2 subtracted from twice the number. Find the number.
Answer by rfer(16322)

3x+3/4=2x-1/2
x=-1/2-3/4
x=-2/4-3/4
x=-5/4
x=-1 1/4